# Getting to the Bottom of Serverless Billing

### v1

@@ 4.1. Overallocation on Public Serverless Platforms
We deploy the same single-threaded, compute-bound serverless function (PyAES from Functionbench (Kim and Lee, 2019)) on AWS Lambda under memory sizes ranging from 128 MB (minimum size) to 1,769 MB, and on GCP (first generation) under CPU configurations ranging from 0.08 (minimum size) to 1 vCPUs. AWS Lambda allocates vCPUs proportionately to the configured memory size, with 1,769 MB equivalent to 1 vCPU (web, 2025k; Zhang et al., 2021), while GCP provides a fine-grained CPU control knob with a 0.01 vCPUs increment (web, 2025l). Figure 10 shows the execution duration reported by the serverless platform with 900,000 samples in total under different CPU configurations. We make two main observations based on these real-world execution logs.
First, a single-threaded, CPU-bound workload like PyAES with a fractional vCPU allocation should experience a slowdown of $\frac{1}{{{\rm{vCPUFraction}}}}$ that follows reciprocal scaling (i.e., half the core allocation, double the execution time). However, the empirical average (solid blue line) is consistently less than the expected average duration (dashed red line) on AWS and GCP (except for a few of the smallest vCPU allocations on GCP). The expected average and expected 5th percentile shown in the figure are based on measurements at full vCPU allocation, scaled proportionally for smaller resource allocations. In other words, a function can ask for, say, half the resources, but be less than twice as slow. In cost terms, this means that users may be charged less than expected under the current wall time billing models presented in Equation 1.
@@
The performance patterns shown in Figure 10 give us clues into what might be going on. Reducing the resource allocation of the AWS Lambda function from 1 vCPU at first does not affect the performance of the function, but suddenly there are increases at slightly above 1400 MB, 700 MB, 470 MB, 350 MB, 280 MB, and so on. These follow a scaled harmonic sequence: ~1400×{$1400\times\{1,\frac{1}{2},\frac{1}{3},\frac{1}{4},\frac{1}{5},...\}$. This discrete $\frac{1}{n}$ sequence suggests the presence of a quantization effect, rather than the continuous proportional allocation ($\frac{1}{x}$) initially expected. Namely, the function is sometimes given more than it is supposed to receive, since the underlying CPU allocation units are quantized, causing jumps on the performance curve. As an analogy, if you want 2 kg of sugar and it is sold in 1 kg packs, the seller gives you two packs. However, if you ask for 1.5 kg, the seller would still need to give you 2 packs, leaving you with an extra 0.5 kg (i.e., overallocation). We observe the same quantization-based overallocation on major serverless platforms.
@@ 4.2. Quantized OS Scheduling
By default, the Linux kernel leverages the Completely Fair Scheduler (CFS) or the Earliest Eligible Virtual Deadline First (EEVDF) scheduler (default scheduler since Linux kernel 6.8) to allocate resources in a fair or latency-sensitive manner (Corbet, 2023). The scheduler generally provides each runnable process with a baseline allocation of resources (e.g., CPU time slice), ensuring that each runnable task can get at least one opportunity to execute on the processor. It also incorporates mechanisms like CPU Bandwidth Control (Turner et al., 2010) and Cgroups (web, 2015b) to impose resource limits and provide resource isolation. Such mechanisms have become the foundation of resource isolation and allocation in the sandboxing solutions widely deployed in serverless, such as containers (web, 2025m; Li et al., 2022), microVMs (Agache et al., 2020), and Wasm (Shillaker and Pietzuch, 2020). Our observations in §4.1 are the results of the existing allocation slices in the OS scheduler and Cgroups, which seem to be coarse-grained for increasingly short serverless functions, causing issues with cost fairness and performance variability.
The OS maintains a kernel data structure (cfs_bandwidth) for bandwidth control of each cgroup (task_group), which includes information such as the enforcement period (CFS period), runtime quota (CFS quota) within each period, remaining runtime available for use (global runtime pool) protected by a spinlock, as well as the throttled run queue (web, 2024e). Note that the newer Linux kernels with the EEVDF scheduler use a similar interface and kernel data structure for CPU bandwidth control as CFS. Therefore, the CFS period and quota we discuss in this section also apply to kernels with the EEVDF scheduler. For the rest of the section, we refer to them as (CPU bandwidth control) period and quota.
A high-resolution timer (hrtimer) is registered with a callback (web, 2024c) to refill the global pool with the quota once per period. Each logical CPU core within the cgroup has a local pool of available runtime for per-CPU-basis runtime accounting. During runtime accounting (e.g., at scheduler ticks or context switches), the consumed runtime is subtracted from the local pool for processes running on a core within the cgroup. When the local pool runs out of runtime, it attempts to acquire more (the smaller of sched_cfs_bandwidth_slice (web, 2015a) or remaining runtime) from the global pool. If both the global and local pools are exhausted, processes on the core are throttled and moved to the throttled run queue. When the global pool is refilled to have available runtime in the new period (hrtimer callback), the scheduler distributes runtime among throttled run queues and unthrottles them (i.e., marks them as eligible to be scheduled again). Under this schema, the wall clock duration of a CPU-bound process can be calculated as:
@@
Here, $d$ is the execution duration, $T$ is the required CPU time, $P$ is the period, and $Q$ is the quota. The scheduler tries to limit the CPU utilization of tasks under CPU bandwidth control to $Q/P$. Figure 11 shows the execution durations derived by Equation 2 for a CPU-bound workload with a CPU time of 33.1 ms (the average value [^2] The requests that report zero CPU usage are excluded. in Huawei serverless traces (Joosen et al., 2025; web, 2024b)) under different periods from 5 ms to 80 ms and the quotas mapped by varying fractional vCPU allocations. These periods are in the same scale compared to those we found empirically (shown later in §4.3). With longer periods, the quantization effect becomes more pronounced. As periods decrease, the execution duration converges to the ideal execution duration following reciprocal scaling.
@@
Figure 11. Shorter CPU bandwidth control periods improve degradation proportionality for sub-core allocations.
@@
The model above does not account for the fact that the runtime accounting and throttling mechanisms cannot operate with infinite frequency or precision due to excessive overhead (e.g., handling hrtimer interrupts (web, 2024a)) in real-world systems. Since the scheduling tick frequency is usually between 100 and 1,000 Hz (CONFIG_HZ(Pellegrini and Quaglia, 2015; web, 2013), runtime accounting and task group throttling is often delayed, especially with the relatively long scheduler tick frequency (e.g., 250 Hz or less). Therefore, a task may often consume runtime more than the quota within a period (overrun) due to lagged accounting, resulting in a negative runtime in the local pool (Ugedal and Kumar, 2022). In this case, the task may be throttled for one or more periods to wait for the quota refill and pay back the runtime debt. For example, consider a CPU-bound task within cgroup with 1.45 ms quota over 20 ms period (i.e., 0.072 vCPU allocated to AWS Lambda with 128 MB memory) and tick interval of 4 ms (250 Hz). A possible scenario is that it first gets 4 ms CPU time and is throttled for 36 ms (rest of the first period and the whole second period) and becomes eligible to run again in the third period (after 40 ms). Then, the task runs another 4 ms after the quota is refilled, causing overrun again with more debt, and is throttled for 56 ms until 100 ms and so on. This task repeatedly alternates between running for 4 ms and being throttled for long periods (i.e., 36 ms or 56 ms) over multiple periods due to overrun and lagged accounting.
Algorithm 1 Profile Runtime and Throttle
1:$s\leftarrow get\_clock\_monotonic()$$\triangleright$ Get monotonic clock time
2:$n\_throt\leftarrow 0$
3:$THRO\leftarrow[]$$\triangleright$ Array of tuples of throttle detected time and throttle duration
4:$last\_chkpt\leftarrow s$
5:while $true$ do
6:    $now\leftarrow get\_clock\_monotonic()$
7:    if $now-last\_chkpt\geq 500us$ then
8:         $THRO[n\_throt++]\leftarrow(now,now-last\_chkpt)$
9:    end if
10:    $last\_chkpt\leftarrow now$
11:    if $now-s\geq EXEC\_DUR$ then return $THRO$
12:    end if
13:end while
@@
Modern kernels often run with the tickless mechanism, with less frequent scheduling interrupts under light loads (web, 2024f; Siddha et al., 2007). Also, scheduling decisions and runtime accounting do not occur only at scheduler ticks. Events like voluntary context switches or interrupts (e.g., hrtimer) can also trigger accounting, rescheduling, or preemption. This can lead to variations in runtime allocation and throttled duration. Overrun issues marginally impact long tasks as the OS scheduler ensures fairness over time, but can significantly affect short tasks. However, a defining feature of serverless is the short execution for the majority of requests (Shahrad et al., 2020b; Joosen et al., 2025, 2023). Therefore, even without the aforementioned overrun effect, CPU overallocation can still happen if a serverless workload is shorter than the CPU bandwidth control enforcement period. For example, a task that requires 10 ms CPU time running within a Cgroup with a 20 ms period of a 10 ms quota is allowed to consume 100% of the CPU during its brief execution, regardless of the configured limit of 0.5 vCPUs. For relatively long tasks that span multiple periods, such overallocation can still happen within the last period before the task is finished. I/O-bound tasks are usually blocked, usually not using CPU while waiting for I/O (e.g., epoll_wait()). However, when the task resumes after data becomes available, overruns and throttling across periods may occur, though this is less pronounced as the task uses the CPU intermittently, consuming less runtime and triggering fewer throttles. In a word, current OS scheduling granularity seems to be coarse in the context of serverless computing.
@@
Figure 12. We successfully match the local scheduling setting to cloud deployments. The scheduler profiling results (figures (a), (b), and (c)) reveal that the scheduling settings and granularity vary across serverless platforms.
@@ 4.3. Scheduling Granularity of Serverless Platforms
Methodology: Algorithm 1 presents the pseudocode of the scheduler profile function, in which the function runs for a predefined duration (EXEC_DUR) and records the time and value of sudden increases (¿500 us) in monotonic clock time (CLOCK_MONOTONIC) readings. The default minimal preemption granularity for CPU-bound tasks in the kernel is 750 us (web, 2024d), and such time jumps can effectively suggest the occurrence of throttles. We invoke the function with different vCPU configurations, each with 300 invocations. Each function request runs for 10 s, leading to runtime/throttle data collected over 3,000 s of execution span for each configuration. Additionally, to be able to assess the effect of different quotas, periods, and OS schedulers, we use in-house VMs, each with 10 vCPUs (Intel Xeon E5-2673 v4), Linux kernel 6.2 (CFS) or 6.8 (EEVDF scheduler), and the timer frequency of either 250 Hz or 1,000 Hz, to profile the function within containers (runC runtime). We analyze the interval between throttles, the throttle duration, and the consumed CPU time before each throttle by calculating the differences between consecutive events in the recorded data.
Empirical Analysis: Figures 12(d)(a) to (c) present the distribution of throttle intervals, durations, and obtained CPU time (runtime) of the studied settings. AWS Lambda functions have throttle intervals that are multiples of 20 ms, whereas IBM functions show multiples of 10 ms. The interval, duration, and runtime results closely align with local runs with corresponding vCPU allocations, periods of 20 ms (for AWS) and 10 ms (for IBM), and the timer frequency of 250 Hz. Also, the runtime and throttle duration of the AWS function (128 MB, 0.072 vCPUs) and their distributions align with the theoretical analysis discussed in § 4.2. The quantized obtained CPU time of AWS Lambda suggests a coarse scheduling granularity under a lower timer frequency (i.e., 250 Hz). The overrun almost happens every time the task is scheduled. Functions on IBM show similar quantized scheduling patterns. The GCP functions exhibit throttle intervals of 100 ms in most cases, while they have 6.42% - 14.83% of throttle durations shorter than 2 ms, indicating frequent context switches and preemption events even within the CPU bandwidth control quota. Compared to other platforms, the less quantized obtained CPU time (i.e., a smoother curve without distinct step-like jumps as shown in  Figure 12(d)(b)-Mid) indicates finer-grained time slice allocation under a higher timer frequency. Table 3 presents the scheduling parameters obtained by our empirical analysis, which suggest that public cloud providers do not have a unanimous configuration.
@@
Table 3. Scheduling parameters obtained by empirical analysis, which vary across different providers.
Does the new EEVDF scheduler solve the overallocation issue? The EEVDF scheduler has replaced the CFS scheduler in Linux kernel version 6.8, which introduces a virtual deadline mechanism that improves system responsiveness by prioritizing latency-sensitive tasks with shorter time slices (Stoica and Abdel-Wahab, 1995; Corbet, 2023). However, overrun issues still persist under EEVDF because runtime accounting and scheduling granularity remain tied to the timer frequency. As shown in Figure 12(d)(d), when using EEVDF with a 250 Hz timer, the CPU time obtained often exceeds the configured quota, though it is slightly better than CFS with less overrun. Raising the timer frequency to 1000 Hz significantly mitigates the overrun issue. However, even with higher timer frequencies, the fundamental overallocation problem still exists. Whenever required CPU time falls below the quota, overallocation cannot be avoided, regardless of scheduler or timer settings.
Implications: Overrun and overallocation are widespread on public serverless platforms. However, providers can absorb this under-accounted resource usage through currently high invocation fees and coarse billing granularity (rounding up), as discussed in §2.5. For example, a GCP function configured with 0.5 vCPUs and 512 MB memory can potentially consume 100% CPU within 50 ms, but GCP will round its billable wall time up to 100 ms plus a high invocation fee equivalent to 30.19 ms. Also, we tested a user-side exploit on AWS Lambda. We implement an intermittent execution framework and decompose a long function (the video-processing application from SeBS (Copik et al., 2021)) into a sequence of short bursts, each falling within the quota. We could reduce billable memory GB-seconds by $66.7\%$ on average (calculated over 100 data points). However, because AWS charges a fixed invocation fee, our actual bill increased by $76.7\%$. In other words, providers that plan to eliminate invocation fees and coarse billing granularity should account for these overallocation effects.

### v2

@@ 4.1. Overallocation on Public Serverless Platforms
We deploy a single-threaded, compute-bound serverless function (PyAES from Functionbench (Kim and Lee, 2019)) on AWS Lambda under memory sizes ranging from 128 MB (minimum size) to 1,769 MB, and on GCP (first generation is used due to its support for fractional vCPU allocation) under CPU configurations ranging from 0.08 (minimum size) to 1 vCPU. AWS Lambda allocates vCPUs proportionately to the configured memory size, with 1,769 MB equivalent to 1 vCPU (Services, 2025c; Zhang et al., 2021), while GCP provides a fine-grained CPU control knob with a 0.01 vCPUs increment (Cloud, 2025h). Figure 10 shows the execution duration reported by the serverless platform with 900,000 samples in total under different CPU configurations. We make two main observations based on these real-world execution logs.
First, a single-threaded, CPU-bound workload like PyAES with a fractional vCPU allocation should experience a slowdown of $\frac{1}{{{\rm{vCPUFraction}}}}$ that follows reciprocal scaling (i.e., half the core allocation, double the execution time). However, the empirical average (solid blue line) is consistently less than the expected average duration (dashed red line) on AWS and GCP (except for a few of the smallest vCPU allocations on GCP). The expected average and expected 5th percentile shown in the figure are based on measurements at full vCPU allocation, scaled proportionally for smaller resource allocations. In other words, a function can ask for, say, half the resources, but be less than twice as slow. In cost terms, this means that users may be charged less than expected under the current wall-clock time billing models presented in Equation 1.
@@
The performance patterns shown in Figure 10 give us clues into what might be going on. Reducing the resource allocation of the AWS Lambda function from 1 vCPU at first does not affect the performance of the function, but suddenly there are increases at slightly above 1400 MB, 700 MB, 470 MB, 350 MB, 280 MB, and so on. These follow a scaled harmonic sequence: $1400\times\{1,\frac{1}{2},\frac{1}{3},\frac{1}{4},\frac{1}{5},$...\}. This discrete $\frac{1}{n}$ sequence suggests the presence of a quantization effect, rather than the continuous proportional allocation ($\frac{1}{x}$) initially expected. Namely, the function is sometimes given more than it is supposed to receive, since the underlying CPU allocation units are quantized, causing jumps on the performance curve. As an analogy, if you want 2 kg of sugar and it is sold in 1 kg packs, the seller gives you two packs. However, if you ask for 1.5 kg, the seller would still need to give you 2 packs, leaving you with an extra 0.5 kg (i.e., overallocation). We observe the same quantization-based overallocation on major serverless platforms.
@@ 4.2. Quantized OS Scheduling
By default, the Linux kernel leverages the Completely Fair Scheduler (CFS) or the Earliest Eligible Virtual Deadline First (EEVDF) scheduler (default scheduler since Linux kernel 6.8) to allocate resources in a fair or latency-sensitive manner (Corbet, 2023). The scheduler generally provides each runnable process with a baseline allocation of resources (e.g., CPU time slice), ensuring that it receives at least one opportunity to execute on the processor. It also incorporates mechanisms like CPU Bandwidth Control (Turner et al., 2010) and cgroups (Kernel, 2015b) to impose resource limits and provide resource isolation. Such mechanisms have become the foundation of resource isolation and allocation in the sandboxing solutions widely deployed in serverless, such as containers (Authors, 2025; Li et al., 2022), microVMs (Agache et al., 2020), and Wasm (Shillaker and Pietzuch, 2020). Our observations in §4.1 are the results of the existing allocation slices in the OS scheduler and cgroups, which seem to be coarse-grained for increasingly short serverless functions, causing issues with cost fairness and performance variability.
The OS maintains a kernel data structure (cfs_bandwidth) for bandwidth control of each cgroup (task_group), which includes information such as the enforcement period (CFS period), runtime quota (CFS quota) within each period, remaining runtime available for use (global runtime pool) protected by a spinlock, as well as the throttled run queue (Kernel, 2024c). Note that the newer Linux kernels with the EEVDF scheduler use a similar interface and kernel data structure for CPU bandwidth control as CFS. Therefore, the CFS period and quota we discuss in this section also apply to kernels with the EEVDF scheduler. For the rest of the section, we refer to them as (CPU bandwidth control) period and quota.
A high-resolution timer (hrtimer) is registered with a callback (Kernel, 2024a) to refill the global pool with the quota once per period. Each logical CPU core within the cgroup has a local pool of available runtime for per-CPU-basis runtime accounting. During runtime accounting (e.g., at scheduler ticks or context switches), the consumed runtime is subtracted from the local pool for processes running on a core within the cgroup. When the local pool runs out of runtime, it attempts to acquire more (the smaller of sched_cfs_bandwidth_slice (Kernel, 2015a) or remaining runtime) from the global pool. If both the global and local pools are exhausted, processes on the core are throttled and moved to the throttled run queue. When the global pool is refilled to have available runtime in the new period (hrtimer callback), the scheduler distributes runtime among throttled run queues and unthrottles them (i.e., marks them as eligible to be scheduled again). Under this schema, the wall clock duration of a CPU-bound process can be calculated as:
@@
Here, $d$ is the execution duration, $T$ is the required CPU time, $P$ is the period, and $Q$ is the quota. The scheduler tries to limit the CPU utilization of tasks under CPU bandwidth control to $Q/P$. Figure 11 shows the execution durations derived by Equation 2 for a CPU-bound workload with a CPU time of 51.8 ms (the average value [^3] The requests that report zero CPU usage are excluded. in Huawei serverless traces (Joosen et al., 2025; Joosen, 2025)) under different periods from 5 ms to 100 ms and the quotas mapped by varying fractional vCPU allocations. These periods are in the same scale compared to those we found empirically (shown later in §4.3). With longer periods, the quantization effect becomes more pronounced. As periods decrease, the execution duration converges to the ideal execution duration following reciprocal scaling.
@@
Figure 11. Theoretical execution durations under fractional CPU allocations. Shorter CPU bandwidth control periods improve degradation proportionality for sub-core allocations.
@@
The model above does not account for the fact that the runtime accounting and throttling mechanisms cannot operate with infinite frequency or precision due to excessive overhead (e.g., handling hrtimer interrupts (Righi, 2024)) in real-world systems. Since the scheduling tick frequency is usually between 100 and 1,000 Hz (CONFIG_HZ(Pellegrini and Quaglia, 2015; Sealey et al., 2013), runtime accounting and task group throttling is often delayed, especially with the relatively long scheduler tick frequency (e.g., 250 Hz or less). Therefore, a task may often consume runtime more than the quota within a period (overrun) due to lagged accounting, resulting in a negative runtime in the local pool (Ugedal and Kumar, 2022). In this case, the task may be throttled for one or more periods to wait for the quota refill and pay back the runtime debt. For example, consider a CPU-bound task within cgroup with 1.45 ms quota over 20 ms period (i.e., 0.072 vCPU allocated to AWS Lambda with 128 MB memory) and tick interval of 4 ms (250 Hz). A possible scenario is that it first gets 4 ms CPU time and is throttled for 36 ms (rest of the first period and the whole second period) and becomes eligible to run again in the third period (after 40 ms). Then, the task runs another 4 ms after the quota is refilled, causing overrun again with more debt, and is throttled for 56 ms until 100 ms and so on. This task repeatedly alternates between running for 4 ms and being throttled for long periods (i.e., 36 ms or 56 ms) over multiple periods due to overrun and lagged accounting.
@@
Modern kernels often run with the tickless mechanism, with less frequent scheduling interrupts under light loads (Kernel, 2024d; Siddha et al., 2007). Also, scheduling decisions and runtime accounting do not occur only at scheduler ticks. Events like voluntary context switches or interrupts (e.g., hrtimer) can also trigger accounting, rescheduling, or preemption. This can lead to variations in runtime allocation and throttled duration. Overrun issues marginally impact long tasks as the OS scheduler ensures fairness over time, but can significantly affect short tasks. However, a defining feature of serverless is the short execution for the majority of requests (Shahrad et al., 2020b; Joosen et al., 2025, 2023). Therefore, even without the aforementioned overrun effect, CPU overallocation can still happen if a serverless workload is shorter than the CPU bandwidth control enforcement period. For example, a task that requires 10 ms CPU time running within a cgroup with a 20 ms period of a 10 ms quota is allowed to consume 100% of the CPU during its brief execution, regardless of the configured limit of 0.5 vCPUs. For relatively long tasks that span multiple periods, such overallocation can still happen within the last period before the task is finished. I/O-bound tasks are usually blocked, usually not using CPU while waiting for I/O (e.g., epoll_wait()). However, when the task resumes after data becomes available, overruns and throttling across periods may occur, though this is less pronounced as the task uses the CPU intermittently, consuming less runtime and triggering fewer throttles. In a word, (I10) current OS scheduling granularity seems to be coarse in the context of serverless computing.
@@
Figure 12. Distributions of throttle intervals, throttle durations, and obtained CPU times (runtime) under the studied scheduling settings. We successfully match the local scheduling setting to cloud deployments. The scheduler profiling results (figures (a), (b), and (c)) reveal that the scheduling settings and granularity vary across serverless platforms.
@@ 4.3. Scheduling Granularity of Serverless Platforms
Algorithm 1 Profile Runtime and Throttle
[…]
5:while truetrue do
6:  $now\leftarrow get\_clock\_monotonic()$
7:  if $now-last\_chkpt\geq 500$$\mu s$ then
8:    $THRO[n\_throt++]\leftarrow(now,now-last\_chkpt)$
9:  end if
10:  $last\_chkpt\leftarrow now$
11:  if $now-s\geq EXEC\_DUR$ then return THROTHRO
12:  end if
13:end while
@@
Methodology: Algorithm 1 presents the pseudocode of the scheduler profile function, in which the function runs for a predefined duration (EXEC_DUR) and records the time and value of sudden increases (¿500 $\mu s$) in monotonic clock time (CLOCK_MONOTONIC) readings. The default minimal preemption granularity for CPU-bound tasks in the kernel is 750 $\mu s$ (Kernel, 2024b), and such time jumps can effectively suggest the occurrence of throttles. We invoke the function with different vCPU configurations, each with 300 invocations. Each function request runs for 10 s, leading to runtime/throttle data collected over 3,000 s of execution span for each configuration. Additionally, to be able to assess the effect of different quotas, periods, and OS schedulers, we use in-house VMs, each with 10 vCPUs (Intel Xeon E5-2673 v4), Linux kernel 6.2 (CFS) or 6.8 (EEVDF scheduler), and the timer frequency of either 250 Hz or 1,000 Hz, to profile the function within containers (runC runtime). We analyze the interval between throttles, the throttle duration, and the consumed CPU time before each throttle by calculating the differences between consecutive events in the recorded data.
Empirical Analysis: Figures 12(a) to (c) present the distribution of throttle intervals, durations, and obtained CPU time (runtime) of the studied settings. AWS Lambda functions have throttle intervals that are multiples of 20 ms, whereas IBM functions show multiples of 10 ms. The interval, duration, and runtime results closely align with local runs with corresponding vCPU allocations, periods of 20 ms (for AWS) and 10 ms (for IBM), and the timer frequency of 250 Hz. Also, the runtime and throttle duration of the AWS function (128 MB, 0.072 vCPUs) and their distributions align with the theoretical analysis discussed in § 4.2. The quantized obtained CPU time of AWS Lambda suggests a coarse scheduling granularity under a lower timer frequency (i.e., 250 Hz). The overrun almost happens every time the task is scheduled. Functions on IBM show similar quantized scheduling patterns. The GCP functions exhibit throttle intervals of 100 ms in most cases, while they have 6.42% - 14.83% of throttle durations shorter than 2 ms, indicating frequent context switches and preemption events even within the CPU bandwidth control quota. Compared to other platforms, the less quantized obtained CPU time (i.e., a smoother curve without distinct step-like jumps as shown in  Figure 12(b)-Mid) indicates finer-grained time slice allocation under a higher timer frequency. Table 3 presents the scheduling parameters obtained by our empirical analysis, which suggest that public cloud providers do not have a unanimous configuration.
@@
Table 3. Scheduling parameters obtained by empirical analysis (as of 2025-05-15), which vary across different providers.
Does the new EEVDF scheduler solve the overallocation issue? The EEVDF scheduler has replaced the CFS scheduler in Linux kernel version 6.8, which introduces a virtual deadline mechanism that improves system responsiveness by prioritizing latency-sensitive tasks with shorter time slices (Stoica and Abdel-Wahab, 1995; Corbet, 2023). However, overrun issues still persist under EEVDF because runtime accounting and scheduling granularity remain tied to the timer frequency. As shown in Figure 12(d), when using EEVDF with a 250 Hz timer, the CPU time obtained often exceeds the configured quota, though it is slightly better than CFS with less overrun. Raising the timer frequency to 1000 Hz significantly mitigates the overrun issue. However, even with higher timer frequencies, the fundamental overallocation problem still exists. Whenever required CPU time falls below the quota, overallocation cannot be avoided, regardless of scheduler or timer settings.
Implications: Overrun and overallocation are widespread on public serverless platforms. However, providers can absorb this under-accounted resource usage through currently high invocation fees and coarse billing granularity (rounding up), as discussed in §2.5. For example, a GCP function configured with 0.5 vCPUs and 512 MB memory can potentially consume 100% CPU within 50 ms, but GCP will round its billable wall-clock time up to 100 ms plus a high invocation fee equivalent to 30.19 ms. Also, we tested a user-side exploit on AWS Lambda. We implement an intermittent execution framework and decompose a long function (the video-processing application from SeBS (Copik et al., 2021)) into a sequence of short bursts, each falling within the quota. We could reduce billable memory GB-seconds by $66.7\%$ on average (calculated over 100 data points). However, because AWS charges a fixed invocation fee, our actual bill increased by $76.7\%$. In other words, providers that plan to eliminate invocation fees and coarse billing granularity should account for these overallocation effects.
@@
One potential way to address overrun and overallocation within the serverless computing context is to adopt an event-driven quota enforcement mechanism instead of periodic polling mechanisms based on periodic timers/ticks (Tsafrir et al., 2007). For example, one-shot timers that expire upon a function process exhausting its bandwidth control quota may be set to trigger an immediate throttle and reschedule. Also, per-task timers can be set to fire after a short, adaptive time (e.g., depending on the global bandwidth control period, overhead tolerance, accuracy requirements, and predicted task duration) to enforce more frequent and accurate CPU time accounting for short-lived tasks with fractional vCPU allocations. In addition, BPF programs can be attached to the scheduler (e.g., through sched_ext (Kernel, 2025)) to selectively apply fine-grained quota enforcement to shorter functions that are more susceptible to overallocation.
5. Discussions
Relative contributions of each cost-related component: In this work, we chose not to quantify the relative contribution of each cost-related component since such numerical breakdowns are highly dependent on context-specific factors. These factors include workload characteristics (e.g., traffic patterns, execution durations, and resource demands), user configurations (e.g., concurrency settings, provisioned resources (Services, 2025d), and subscription plans (Azure, 2025h)), and provider-specific policies (e.g., ARM CPU and committed use discounts and free tiers (Poccia, 2023; Services, 2025f)), which vary across applications and providers. Therefore, any numerical breakdown would not be broadly applicable. Instead, our approach decomposes the inherent sources of cost inefficiencies and presents a systematic analysis framework across multiple abstraction layers from user-facing billing models to OS scheduling, which enables practitioners to measure and rank cost drivers within their own context.
Actionables for serverless users: Our findings lead to several actionable recommendations for reducing serverless costs. First, users can conduct trace-based analysis to pick an appropriate platform whose cost drivers, such as billing practices (§2 and Table 1), concurrency modes (§3.1), serving architectures (§3.2), keep-alive patterns (§3.3), and scheduling granularity (§4.3 and Table 3), best match their workload. Depending on the cost breakdown, users may consider merging similar functions to lower invocation fees, decomposing functions to better utilize resources, or configuring always-ready instances to avoid cold starts (Baldini et al., 2017; Yu et al., 2020; Mahgoub et al., 2022a; Services, 2025d). Also, users should be wary of serverless concurrency models and tune control knobs for resources and scaling to avoid the dual penalty of slowdowns and higher bills (§3.1). Furthermore, it is a good practice to tune workload resource demands and fractional vCPU allocations to avoid performance jitters due to coarse OS scheduling granularity (i.e., quantization jumps shown in Figure 10). Lastly, serverless users may also consider the possibility of running background tasks during keep-alive periods (§3.3 and Table 2).